OpenStudy (anonymous):

Please Help I Will give Medal and Fan!!! A triangle has base (x-5) units and height (4x-10) units. A. What is the square of the length of the hypotenuse of the triangle? Show your work. B. What is the are of the triangle? C. Using the solution obtained in Part B, explain the Closure Property of Multiplication of Polynomials.

OpenStudy (sidsiddhartha):

a=x-5 b=4x-10 c^2=a^2+b^2 so c will be length of hypotenuse and area of the triangle=1/2*(base)*(height)

OpenStudy (anonymous):

So is a b and c the triangle?

OpenStudy (sidsiddhartha):

yeah the above theorem is known as Pythagorean theorem

OpenStudy (anonymous):

Do u mind putting this into an equation?

OpenStudy (anonymous):

nevermind that.

OpenStudy (anonymous):

So is it 5x-125

OpenStudy (anonymous):

Now I need help with part C